If A 2 -A+I=0, then what is A inverse ?


If A2-A+I=0,  then what is A inverse ?


Grade:11

1 Answers

Askiitians Expert Sahil Arora - IITD
13 years ago

Dear sudeep

As Identity is present in the equation, we can get A-1 simply by multiplying.....

Multiplying both sides by A-1,we get

A-I+A-1=0

==>A-1=I-A

Please feel free to post as many doubts on our discussion forum as you can. If you find any question 
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We 
are all IITians and here to help you in your IIT JEE preparation. 


All the best sudeep !!!




Regards,

Askiitians Experts

Sahil Arora